OpenStudy (anonymous):

how would you solve this type of equation: -12+2v=-3(8v+4). Can somone help me with the steps???

OpenStudy (anonymous):

you know you got to distribute the 3 into the parentheses first

OpenStudy (anonymous):

i know you have to distrubute first but when i try to just solve for v i always end up with zero, for all the problems.... so i think i'm doing something wrong.

OpenStudy (anonymous):

multply -3 with(8v+4) -12+2V=-24v-12 24v+2v=12-12 v=0